OpenStudy (anonymous):

what is the solution of this equation x^2-2x+4=0

myininaya (myininaya):

you should use the quadratic formula or solve by completing the square

myininaya (myininaya):

\[ax^2+bx+c=0 => x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

a=1,b=-2,c=4 just plug into the formula i just gave you

OpenStudy (anonymous):

We'll see that no real solution exists here.

OpenStudy (anonymous):

X^2 - 2X + 4 = 0 , X = [ 2 +/- (-12) ^ 1/2 ] /2 = 1 +/- ( 3) ^ 1/2 i